# Implications of DESI for Dark Matter & Cosmic Birefringence

### v1

@@ 1 Introduction
Measurements of Baryon Acoustic Oscillations (BAO) from the Dark Energy Spectroscopic Instrument (DESI) DESI:2024mwx , combined with type Ia Supernova observations from the Union3 compilation SupernovaCosmologyProject:1998vns and the Dark Energy Survey Year-5 data DES:2024jxu , suggest the possibility that dark energy (DE) may be evolving throughout cosmic history. Notably, the data favor a DE equation of state (EoS) $w<-1$ at high redshifts and $w>-1$ at present, indicating a so-called “phantom crossing” behavior Hu:2004kh ; Guo:2004fq . The most recent DESI DR2 results DESI:2025zgx ; DESI:2025fii appear to favor such a scenario, though the uncertainties are still substantial. Future measurements with improved precision will be crucial to confirm whether this crossing reflects a genuine feature of the dark energy dynamics.
@@
Constructing consistent models that yield an equation-of-state parameter $w<-1$ has proven to be theoretically challenging. Many existing approaches, such as phantom models Caldwell:2003vq , invoke ghost-like degrees of freedom, which often lead to instabilities and other theoretical pathologies Carroll:2003st ; Cline:2003gs ; Hsu:2004vr . While scalar-tensor theories with non-minimal couplings to gravity, such as Brans–Dicke theory, can in principle reproduce $w<-1$ behavior Boisseau:2000pr ; Torres:2002pe ; Capozziello:2002rd ; Carroll:2004hc ; Liddle:2003as , stringent solar-system constraints render the Brans–Dicke scalar effectively inert, thereby forcing $w\approx-1$. A well-studied class of models is quintessence Copeland:2006wr ; PhysRevD.37.3406 ; Zlatev:1998tr . In quintessence models, the evolution of the scalar field generally falls into two distinct categories: freezing models, where the field slows down over time and asymptotically approaches $w\to-1$, and thawing models, where the field remains initially frozen near $w\simeq-1$ and begins to evolve only recently. Although a variety of potentials, such as exponential, power-law, and others, have been extensively studied, they universally predict an equation-of-state parameter in the range $-1\leq w<1/3$ Tsujikawa:2013fta ; Bhattacharya:2024hep ; Luu:2025fgw . This range supports cosmic acceleration but does not extend into the phantom regime, making it challenging for such models to accommodate the behavior suggested by the DESI results Wolf:2024eph ; Ramadan:2024kmn ; Bhattacharya:2024kxp . Alternative proposals for achieving $w<-1$ include brane-world models Sahni:2002dx , quantum corrections Onemli:2002hr , couplings between quintessence and moduli fields Chung:2002xj , and photon–axion conversion mechanisms Csaki:2004ha . Moreover, the observed phantom-like behavior may also arise from a non-minimal coupling of DE to gravity Carroll:2004hc , or through interactions between the DE and dark matter (DM) Huey:2004qv ; Das:2005yj ; Chakraborty:2025syu . Very recently, such models have also been tested using the gravitational-wave observations from LIGO and VIRGO Yang:2019vni ; Bachega:2019fki ; Li:2019ajo . Models of interacting DE and DM have also been shown to alleviate the $S_{8}$ tension, which is an apparent discrepancy between the amplitude of matter clustering inferred from the cosmic microwave background (CMB) and that observed at late times in weak lensing surveys Poulin:2022sgp ; Khoury:2025txd . Models featuring DE–DM interactions could offer an appealing twofold resolution: they not only allow $w<-1$, and help to alleviate the $S_{8}$ tension, but may also hint at the origin of DM. Interacting DM-DE models have been studied in various contexts, see, for example, Refs. Damour:1990tw ; Casas:1991ky ; Bean:2001ys ; Comelli:2003cv ; Franca:2003zg ; Chimento:2003iea ; Amendola:1999er ; Amendola:2000uh ; Amendola:2003eq ; Olivares:2005tb ; Gubser:2004uf ; Brax:2004qh ; Copeland:2006wr ; Bolotin:2013jpa , however, they lack a consistent picture of particle DM production.
@@
Motivated by these considerations, the present work aims to simultaneously address two key questions: (a) how to construct a DE model that yields an equation-of-state parameter $w<-1$ at low redshifts, and (b) how such a model can concurrently account for DM genesis within an interacting DE–DM framework, while also addressing the non-zero ICB angle. To this end, we build upon the model proposed in Ref. Das:2005yj , where a canonical scalar field, serving as the quintessence DE component, interacts with a fermionic DM. We extend this setup by introducing a dimension-5 operator coupling the scalar field to the electromagnetic (EM) field strength tensor. We consider two representative forms for the DE scalar potential: a polynomial potential and an axion-like potential. We give an ultraviolet (UV) completion for the model proposed in Ref. Khoury:2025txd and demonstrate how the EM coupling can be induced, which explains the observed ICB angle. Interestingly, the dynamics of the DS lead to very different DM mass scales in the two cases: while the polynomial potential supports $\gtrsim\mathcal{O}(10^{10})$ GeV DM (produced via freeze-in), the axionic scenario favors a GeV-TeV DM candidate (produced from DS freeze-out). We also outline the possibility of asymmetric DM production in the latter case, which requires additional extension of the model parameters, but could potentially give rise to DM self-interaction strength desirable from the explanation of diversity of rotation curves of spiral galaxies Roberts:2024uyw . In both frameworks, the viable parameter space is found to be consistent with the observed DM relic abundance, recent DESI data, and the measured ICB angle.
@@ 2 The framework
where $V(\varphi)$ is the DE potential, $\alpha_{\rm em}$ is the EM fine-structure constant, $F^{\mu\nu}$ is the EM field strength tensor, $\tilde{F}^{\mu\nu}=(1/2)\epsilon^{\mu\nu\rho\sigma}F_{\rho\sigma}$ is its dual, and $f$ is the characteristic scale associated with $\varphi$. For instance, if $\varphi$ is a quintessence axion, $f$ plays the role of its decay constant. Notice we have introduced the anomalous EM coupling of $\varphi$ having Wilson coefficient $c_{\gamma}$ that could explain the ICB Lin:2013sca . The source of DM-DE interaction is the $\varphi$ dependent DM mass, denoted as $m(\varphi)$ in Eq. (1). If $\varphi$ is a pseudo-scalar like an axion, associated with a dark QCD sector, then the anomalous coupling with $F^{\mu\nu}$ is natural, and can be induced if the hypercharge current is anomalous under the non-linearly realized shift symmetry of the axion. In this case, the scalar-like coupling with $\psi$ can appear as a result of finite density DM effects on the dark QCD condensate, provided $\psi$ is identified with the stable dark baryon DM Khoury:2025txd . If, on the other hand, $\varphi$ is a scalar, the required EM coupling can be generated if there are parity-violating interactions in the DS such that parity-odd and even states can mix. If there is no symmetry protecting the DE mass and potential shape, interactions with the DM can significantly raise the DE mass and destroy its potential shape, especially when there is a large hierarchy between the DE and DM masses DAmico:2016jbm . However, if there is a remnant symmetry, such as a non-linearly realized shift symmetry of a broken scale/chiral invariance, that would protect the DE mass and potential shape. The scale invariance is suitable for a scalar (dilaton) like DE candidate, while the remnant shift symmetry can protect the pseudoscalar (axion) DE mass and potential against radiative instability.
@@ 2.1 Evolution of the DE and effective equation of state
where $M_{P}$ is the reduced Planck mass, $a$ denotes the scale factor, the label ‘00’ indicates the present-day value of the corresponding quantities, $\rho_{\rm DM}$, $\rho_{\rm B}$, and $\rho_{\rm rad}$ stand for the energy density stored in DM, baryons, and radiation component, while the DE energy density $\rho_{\varphi}$ is given as
@@
where $\theta^{\prime}(a)\equiv d\theta/da$. One can consistently solve Eqs. (5), (8) for some given $V(\theta),\ m(\theta)$ for a specific choice of $\tilde{f},\theta_{0}$ by adjusting the initial conditions, and obtain the effective EoS for DE as a function of the scale factor or redshift utilizing (14), and fit it with the DESI results. We will consider some concrete examples, and elaborate on the behavior of $w_{\rm eff}$. Note from Eq. (14) that, today $w_{\rm eff}$ approaches $-1$ if $\tilde{f}\ll 1$.
Cosmic birefringence
@@ 2.2 Polynomial potential
BP1 0.2 0.08 1 2.59 0.39 $7.7-12.4$
BP2 $5\times 10^{-3}$ $5\times 10^{-3}$ $10^{-2}$ 2.74 1.86 $1.6-2.6$
BP3 $3\times 10^{-3}$ $8\times 10^{-3}$ $10^{-3}$ $2.74$ $0.97$ $3-5$
@@ 2.3 Axion type potential
Consider an $SU(N_{\rm D})$ dark QCD sector with two light quarks denoted as $\mathbf{u},\mathbf{d}$. A global $U(1)_{\rm PQ}$ Peccei-Quinn (PQ) symmetry is anomalous under $SU(N_{\rm D})$ and is broken spontaneously at a scale $f$, which results in a pseudo-Goldstone boson axion $\varphi$, that transforms as: $\varphi\to\varphi+\mathbb{T}f$ under the $U(1)_{\rm PQ}$ with the transformation parameter $\mathbb{T}$. The field $\varphi$ has the following coupling with dark gluon field strength $G_{\rm D}$
${\cal L}_{\rm D}=\frac{g_{\rm D}^{2}}{32\pi^{2}}\frac{\varphi}{f}{\rm Tr}\left[G_{\rm D}^{\mu\nu}\widetilde{G}_{{\rm D};\mu\nu}\right]\ ,$ (18)
where $g_{\rm D}$ stands for the $SU(N_{\rm D})$ coupling, and $\widetilde{G}_{\rm D}^{\mu\nu}=\frac{1}{2}\epsilon^{\mu\nu\rho\sigma}G_{\rm D;\rho\sigma}$. This can be UV completed, for example, by introducing a pair of left-handed chiral fermions $\{Q,\bar{Q}\}$ having representation $(N_{\rm D},1,1,C_{Y})$, $(\bar{N}_{\rm D},1,1,-C_{Y})$ under the extended SM gauge symmetry $SU(N_{\rm D})\otimes SU(3)_{c}\otimes SU(2)_{L}\otimes U(1)_{Y}$, respectively. Here, the overbar does not denote any conjugation. The global PQ charges of $Q,\bar{Q}$ are $1$, 00, respectively. A SM singlet scalar $\Phi$ with PQ charge $1$ obtains a vacuum expectation value (VEV) at a scale $f$, spontaneously breaking $U(1)_{\rm PQ}$ and giving masses to $Q,\bar{Q}$ via its Yukawa coupling $y_{Q}\Phi^{*}\bar{Q}Q+{\rm h.c.}$ The axion $\varphi$ appears as the axial mode of $\Phi$ after it obtains a VEV. The field content of the model is summarized in Tab. 2. Note that, $Q$, $\bar{Q}$ can decay to $\mathbf{u,d}$ and SM particles in the early Universe if $C_{Y}$ is chosen appropriately. For example, the decay can proceed via effective operators like $(\langle\Phi\rangle^{*}/\Lambda_{d}^{3})\bar{\mathbf{u}}_{R}Qe_{R}e_{R}$ , $(1/\Lambda_{d}^{2})\bar{Q}\mathbf{u}_{L}e_{R}^{c}e_{R}^{c}$ if $C_{Y}=2$, while $c$ denotes charge conjugation. As these operators do not violate $U(1)_{\rm PQ}$, they do not contribute to the axion potential. We find that the decay happens much before BBN for any $\Lambda_{d}<M_{P}$ as $m_{Q}\sim f$ for order one $y_{Q}$, thus posing no threat to the standard cosmology. We have also checked that there is no Landau pole for $U(1)_{Y}$ coupling below $M_{P}$ for $N_{\rm D}C_{Y}^{2}=c_{\gamma}\lesssim 18$ for $Q,\bar{Q}$ mass above $10^{12}$ GeV. As $Q,\bar{Q}$ mass is $\simeq y_{Q}f$, this is easily satisfied for order one $y_{Q}$. Note that the light quarks $m_{\mathbf{q}}$ do not cause any Landau pole as they do not possess any hypercharge.
@@
and $\Upsilon\simeq 1$ for $m_{\mathbf{u}}\simeq m_{\mathbf{d}}$ such that for present-day values of $\theta\to 0$, the DM mass is $m_{\rm DM}\simeq m_{\psi}$.
In Fig. 2 we demonstrate the evolution of the axion field and the resulting equation of motion following Ref. Khoury:2025txd . The field naturally finds itself at $\theta=\pi$ irrespective of the initial condition, and evolves to $\theta=0$ in the late Universe, once the DM density has dropped below $n_{\psi}^{c}$. This is a very appealing feature, as it naturally explains why the axion finds itself near the hilltop of the late-time effective potential to explain the DE today. In conventional quintessence axion models, this initial condition is chosen by hand. We obtain the resulting field excursion $\Delta\theta$ from the solution to the EoM for axion, and evaluate the required $c_{\gamma}$ to explain the ICB reported, as shown in Tab. 3. As $c_{\gamma}=N_{\rm D}\,C_{Y}^{2}$, it is relatively easier to accommodate the required charge with ${\cal O}(1)$ $C_{Y}$ for an appropriate number of dark colors $N_{\rm D}$. The chosen benchmark can fit the DESI results, while when extrapolated to higher redshift, the EoS approaches 00 as consistent with the generic interacting DE-DM framework outlined in Sec. 2.

### v2

@@ 1 Introduction
Measurements of Baryon Acoustic Oscillations (BAO) from the Dark Energy Spectroscopic Instrument (DESI) DESI:2024mwx , combined with type Ia Supernova observations from the Union3 compilation SupernovaCosmologyProject:1998vns and the Dark Energy Survey Year-5 data DES:2024jxu , suggest the possibility that dark energy (DE) may be evolving throughout cosmic history. Notably, the data favor a DE equation of state (EoS) w<1w<-1 at high redshifts and w>1w>-1 at present, indicating a so-called “phantom crossing” behavior Hu:2004kh ; Guo:2004fq . The most recent DESI DR2 results DESI:2025zgx ; DESI:2025fii appear to favor such a scenario, though the uncertainties are still substantial. Future measurements with improved precision will be crucial to confirm whether this crossing reflects a genuine feature of the dark energy dynamics.
@@
Constructing consistent models that yield an equation-of-state parameter w<1w<-1 has proven to be theoretically challenging. Many existing approaches, such as phantom models Caldwell:2003vq ; Elizalde:2004mq , invoke ghost-like degrees of freedom, which often lead to instabilities and other theoretical pathologies Carroll:2003st ; Cline:2003gs ; Hsu:2004vr . While scalar-tensor theories with non-minimal couplings to gravity, such as Brans–Dicke theory, can in principle reproduce w<1w<-1 behavior Boisseau:2000pr ; Torres:2002pe ; Capozziello:2002rd ; Carroll:2004hc ; Liddle:2003as , stringent solar-system constraints render the Brans–Dicke scalar effectively inert, thereby forcing $w\approx-1$. A well-studied class of models is quintessence Copeland:2006wr ; PhysRevD.37.3406 ; Zlatev:1998tr ; Bamba:2012cp . In quintessence models, the evolution of the scalar field generally falls into two distinct categories: freezing models, where the field slows down over time and asymptotically approaches $w\to-1$, and thawing models, where the field remains initially frozen near $w\simeq-1$ and begins to evolve only recently. Although a variety of potentials, such as exponential, power-law, and others, have been extensively studied, they universally predict an equation-of-state parameter in the range $1\leq w<1/3$ Tsujikawa:2013fta ; Bhattacharya:2024hep ; Luu:2025fgw . This range supports cosmic acceleration but does not extend into the phantom regime, making it challenging for such models to accommodate the behavior suggested by the DESI results Wolf:2024eph ; Ramadan:2024kmn ; Bhattacharya:2024kxp . Alternative proposals for achieving w<1w<-1 include brane-world models Sahni:2002dx , quantum corrections Onemli:2002hr , pixelated Universe Heckman:2024apk , couplings between quintessence and moduli fields Chung:2002xj , non-minimally coupled quintessence field Wolf:2024stt ; Wolf:2025jed , and photon–axion conversion mechanisms Csaki:2004ha . Moreover, the observed phantom-like behavior may also arise from a non-minimal coupling of DE to gravity Carroll:2004hc , or through interactions between the DE and dark matter (DM) Huey:2004qv ; Das:2005yj ; Chakraborty:2025syu  [^1] In Berbig:2024aee another possibility has been discussed, where a quintessence field possesses a non-zero initial velocity.. Very recently, such models have also been tested using the gravitational-wave observations from LIGO and VIRGO Yang:2019vni ; Bachega:2019fki ; Li:2019ajo . Models of interacting DE and DM have also been shown to alleviate the $S_{8}$ tension, which is an apparent discrepancy between the amplitude of matter clustering inferred from the cosmic microwave background (CMB) and that observed at late times in weak lensing surveys Poulin:2022sgp ; Khoury:2025txd . Models featuring DE–DM interactions could offer an appealing twofold resolution: they not only allow w<1w<-1, and help to alleviate the $S_{8}$ tension, but may also hint at the origin of DM. Interacting DM-DE models have been studied in various contexts [^2] In related models of DE conformally coupled with matter, dubbed as “chameleon dark energy”, the interaction with local matter overdensities can effectively raise the local expansion rate, potentially resolving the Hubble tension without conflicting with large-scale observations Cai:2021wgv (see also Ref. Karwal:2021vpk )., see, for example, Refs. Damour:1990tw ; Casas:1991ky ; Bean:2001ys ; Comelli:2003cv ; Franca:2003zg ; Chimento:2003iea ; Amendola:1999er ; Amendola:2000uh ; Amendola:2003eq ; Olivares:2005tb ; Gubser:2004uf ; Brax:2004qh ; Copeland:2006wr ; Bolotin:2013jpa ; Andriot:2025los ; Li:2024qso , however, they lack a consistent picture of particle DM production.
@@
Motivated by these considerations, the present work aims to simultaneously address two key questions: (a) how to construct a DE model that yields an equation-of-state parameter w<1w<-1 at low redshifts, and (b) how such a model can concurrently account for DM genesis within an interacting DE–DM framework, while also addressing the non-zero ICB angle. To this end, we build upon the model proposed in Ref. Das:2005yj , where a canonical scalar field, serving as the quintessence DE component, interacts with a fermionic DM. We extend this setup by introducing a dimension-5 operator coupling the scalar field to the electromagnetic (EM) field strength tensor. We consider two representative forms for the DE scalar potential: a polynomial potential and an axion-like potential. We give an ultraviolet (UV) completion for the model proposed in Ref. Khoury:2025txd and demonstrate how the EM coupling can be induced, which explains the observed ICB angle. Interestingly, the dynamics of the DS lead to very different DM mass scales in the two cases: while the polynomial potential supports $\gtrsim\mathcal{O}(10^{10})$ GeV DM (produced via freeze-in), the axionic scenario favors a GeV-TeV DM candidate (produced from DS freeze-out). We also outline the possibility of asymmetric DM production in the latter case, which requires additional extension of the model parameters, but could potentially give rise to DM self-interaction strength desirable from the explanation of diversity of rotation curves of spiral galaxies Roberts:2024uyw . In both frameworks, the viable parameter space is found to be consistent with the observed DM relic abundance, recent DESI data, and the measured ICB angle.
@@ 2 The framework
where $V(\varphi)$ is the DE potential, $\alpha_{\rm em}$ is the EM fine-structure constant, $F^{\mu\nu}$ is the EM field strength tensor, $\tilde{F}^{\mu\nu}=(1/2)\epsilon^{\mu\nu\rho\sigma}F_{\rho\sigma}$ is its dual, and $f$ is the characteristic scale associated with $\varphi$. For instance, if $\varphi$ is a quintessence axion, $f$ plays the role of its decay constant. Notice we have introduced the anomalous EM coupling of $\varphi$ having Wilson coefficient $c_{\gamma}$ that could explain the ICB Lin:2013sca . The source of DM-DE interaction [^3] If there is no symmetry protecting the DE mass and potential shape, interactions with the DM can significantly raise the DE mass and destroy its potential shape, especially when there is a large hierarchy between the DE and DM masses DAmico:2016jbm . However, if there is a remnant symmetry, such as a non-linearly realized shift symmetry of a broken scale/chiral invariance, that would protect the DE mass and potential shape. The scale invariance is suitable for a scalar (dilaton) like DE candidate (see for e.g., Ref. Brax:2004ym ), while the remnant shift symmetry can protect the pseudoscalar (axion) DE mass and potential against radiative instability. is the $\varphi$ dependent DM mass, denoted as $m(\varphi)$ in Eq. (1). If $\varphi$ is a pseudo-scalar like an axion, associated with a dark QCD sector, then the anomalous coupling with $F^{\mu\nu}$ is natural, and can be induced if the hypercharge current is anomalous under the non-linearly realized shift symmetry of the axion. In this case, the scalar-like coupling with $\psi$ can appear as a result of finite density DM effects on the dark QCD condensate, provided $\psi$ is identified with the stable dark baryon DM Khoury:2025txd . If, on the other hand, $\varphi$ is a scalar, the required EM coupling can be generated if there are parity-violating interactions in the DS such that parity-odd and even states can mix.
@@ 2.1 Evolution of the DE and effective equation of state
where $M_{P}$ is the reduced Planck mass, $a$ denotes the scale factor, the label ‘0’ indicates the present-day value of the corresponding quantities, $\rho_{\rm DM}$, $\rho_{\rm B}$, and $\rho_{\rm rad}$ stand for the energy density stored in DM, baryons, and radiation component, while the DE energy density $\rho_{\varphi}$ is given as
@@
where $\theta^{\prime}(a)\equiv d\theta/da$. One can consistently solve Eqs. (5), (8) for some given $V(\theta),\ m(\theta)$ for a specific choice of $\tilde{f},\theta_{0}$ by adjusting the initial conditions, and obtain the effective EoS for DE as a function of the scale factor or redshift utilizing (14), and fit it with the DESI results. We will consider some concrete examples, and elaborate on the behavior of $w_{\rm eff}$. Note from Eq. (14) that, today $w_{\rm eff}$ approaches 1-1 if $\tilde{f}\ll 1$.
Isotropic cosmic birefringence
@@ 2.2 Polynomial potential
BP1 0.2 0.08 1 2.59 0.39 8128-12
BP2 $5\times 10^{-3}$ $5\times 10^{-3}$ $10^{-2}$ 2.74 1.86 232-3
BP3 $3\times 10^{-3}$ $8\times 10^{-3}$ $10^{-3}$ $2.74$ $0.97$ 353-5
@@ 2.3 Axion type potential
Consider an $SU(N_{\rm D})$ dark QCD sector with two light quarks denoted as $\mathbf{u},\mathbf{d}$. A global $U(1)_{\rm PQ}$ Peccei-Quinn (PQ) symmetry is anomalous under $SU(N_{\rm D})$ and is broken spontaneously at a scale $f$, which results in a pseudo-Goldstone boson axion $\varphi$, that transforms as: $\varphi\to\varphi+\mathbb{T}f$ under the $U(1)_{\rm PQ}$ with the transformation parameter $\mathbb{T}$. The field $\varphi$ has the following coupling with dark gluon field strength $G_{\rm D;\mu\nu}^{a}$
${\cal L}_{\rm D}=\frac{g_{\rm D}^{2}}{32\pi^{2}}\frac{\varphi}{f}G_{\rm D}^{a\,\mu\nu}\widetilde{G}^{a}_{{\rm D};\mu\nu}\ ,$ (18)
where $g_{\rm D}$ stands for the $SU(N_{\rm D})$ coupling, and $\widetilde{G}^{a}_{\rm D;\mu\nu}=\frac{1}{2}\epsilon_{\mu\nu\rho\sigma}G_{\rm D}^{a\,\rho\sigma}$. This can be UV completed, for example, by introducing a pair of left-handed chiral fermions $\{Q,\bar{Q}\}$ having representation $(N_{\rm D},1,1,C_{Y})$, $(\bar{N}_{\rm D},1,1,-C_{Y})$ under the extended SM gauge symmetry $SU(N_{\rm D})\otimes SU(3)_{c}\otimes SU(2)_{L}\otimes U(1)_{Y}$, respectively. The bar in $\bar{Q}$ does not denote any charge conjugation. The global PQ charges of $Q,\bar{Q}$ are $1$, 0, respectively. A SM singlet scalar $\Phi$ with PQ charge $1$ obtains a vacuum expectation value (VEV) at a scale $f$, spontaneously breaking $U(1)_{\rm PQ}$ and giving masses to $Q,\bar{Q}$ via its Yukawa coupling $y_{Q}\Phi^{*}\bar{Q}Q+{\rm h.c.}$ The axion $\varphi$ appears as the axial mode of $\Phi$ after it obtains a VEV. The field content of the model is summarized in Tab. 2. Note that, $Q$, $\bar{Q}$ can decay to $\mathbf{u,d}$ and SM particles in the early Universe if $C_{Y}$ is chosen appropriately. For example, the decay can proceed via effective operators like $(\langle\Phi\rangle^{*}/\Lambda_{d}^{3})\bar{\mathbf{u}}_{R}Qe_{R}e_{R}$ , $(1/\Lambda_{d}^{2})\bar{Q}\mathbf{u}_{L}e_{R}^{c}e_{R}^{c}$ if $C_{Y}=2$, while $c$ denotes charge conjugation. Another choice of operators could be $(\langle\Phi\rangle^{*}/\Lambda_{l}^{3})\bar{\mathbf{u}}_{R}Qe_{L}\nu_{L}$ , $(1/\Lambda_{l}^{2})\bar{Q}\mathbf{u}_{L}e_{L}^{c}\nu_{L}^{c}$ if $C_{Y}=1$. As these operators do not violate $U(1)_{\rm PQ}$, they do not contribute to the axion potential. We find that the decay happens much before BBN for any $\Lambda_{d}<M_{P}$ as $m_{Q}\sim f$ for order one $y_{Q}$, thus posing no threat to the standard cosmology. We have also checked that there is no Landau pole for $U(1)_{Y}$ coupling below $M_{P}$ for $N_{\rm D}C_{Y}^{2}=c_{\gamma}\lesssim 18$ for $Q,\bar{Q}$ mass above $10^{12}$ GeV. As $Q,\bar{Q}$ mass is $\simeq y_{Q}f$, this is easily satisfied for order one $y_{Q}$. Note that the light quarks $\mathbf{q}=\mathbf{u},\mathbf{d}$} do not cause any Landau pole as they do not possess any hypercharge.
@@
and $\Upsilon\simeq 1$ for $m_{\mathbf{u}}\simeq m_{\mathbf{d}}$ such that for present-day values of $\theta\to 0$, the DM mass is $m_{\rm DM}\simeq m_{\psi}$, as $\sigma_{\psi}/m_{\psi}\simeq{\cal O}(10^{-2})$ as shown in Tab. 3.
In Fig. 2 we demonstrate the evolution of the axion field and the resulting equation of motion following Ref. Khoury:2025txd . The field naturally finds itself at $\theta=\pi$ irrespective of the initial condition, and evolves to $\theta=0$ in the late Universe, once the DM density has dropped below $n_{\psi}^{c}$. This is a very appealing feature, as it naturally explains why the axion finds itself near the hilltop of the late-time effective potential to explain the DE today. In conventional quintessence axion models, this initial condition is chosen by hand. We obtain the resulting field excursion $\Delta\theta$ from the solution to the EoM for axion, and evaluate the required $c_{\gamma}$ to explain the ICB reported, as shown in Tab. 3. As $c_{\gamma}=N_{\rm D}\,C_{Y}^{2}$, it is relatively easier to accommodate the required charge with ${\cal O}(1)$ $C_{Y}$ for an appropriate number of dark colors $N_{\rm D}$. To be quantitative, for the BPa, one can choose $N_{\rm D}=3$, and $C_{Y}=2$, while for BPb, $N_{\rm D}=2$, $C_{Y}=1$ is sufficient to obtain the required $c_{\gamma}$ as shown in Tab. 3. The chosen benchmark can fit the DESI results, while when extrapolated to higher redshift, the EoS approaches 0 as consistent with the generic interacting DE-DM framework outlined in Sec. 2.